OpenStudy (anonymous):

The following function defines a recursive sequence. f(0) = -3 f(1) = 6 f(n) = -2•f(n -1) - f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -3, 6, -9, -12, … -3, 20, -95, 480, … -3, 6, -9, 12, … -3, -20, -95, -480, …

OpenStudy (anonymous):

I have never been thought this for some reason I am looking for someone to tell me how to do it.

OpenStudy (anonymous):

taught*

hartnn (hartnn):

ok, f(0) = -3 , first term = -3 f(1) = 6, 2nd term = 6 so, as of now our sequence is -3,6 to find nect term, we will use f(n) = -2•f(n -1) - f(n - 2) put n= 2 here, what do u get ?

OpenStudy (anonymous):

Im sorry, but I lost you at f(n) = -2•f(n -1) - f(n - 2)

hartnn (hartnn):

thats what is already given. you just replace "n" by "2" there, to get f(2)

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

So it would be like f(2) = -2 * f(n-1) - f(n - 2) to keep going correct?

hartnn (hartnn):

you need to replace n by 2 on right side too!

OpenStudy (anonymous):

Oh so.. f(2) = -2 * f(2 - 1) - f(2 - 2)?

hartnn (hartnn):

yes! f(2) = -2 f(1) - f(0) and you already know f(1) =6, f(0) = -3 just plug in values!

OpenStudy (anonymous):

Oh ok so the answer would be the first one right?

hartnn (hartnn):

you need to find f(3) too! f(3) = - 2f(2) - f(1) = - 2(-9) -6 = 18-6 = 12

hartnn (hartnn):

so no, its not the first one

OpenStudy (anonymous):

Oh ok since 18 - 6 is positive 12 not negative. I messed up on that last part I thought it was 6- 18. Thank you I understand now!

hartnn (hartnn):

welcome ^_^